Twice the greatest of three consecutive even integers is 4 more than the smallest of the three numbers. Find the numbers.
Answer by mananth(16949)

let the numbers be x, x+2 , x+4
2(x+4) = x+4
2x+8=x+4
2x-x=4-8
x=-4
..
the numbers are -4,-2,0